OpenStudy (anonymous):

Simplify the following expression: (5x^2 + 3x + 4) - (2x^2 + 5x - 1). If the final answer is written in the form Ax^2 + Bx + C, what is the value of A?

OpenStudy (hoblos):

(5x^2 + 3x + 4) - (2x^2 + 5x - 1) = 5x^2 + 3x + 4 -2x^2 -5x +1 = 3x^2 -2x +5 A=3

OpenStudy (anonymous):

Ur amazing!!! @hoblos

OpenStudy (hoblos):

:D:D thanks